Session: Oct 2011 (session 3)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Franco Fiocchi at Sun Nov 06 00:08:29 2011) Failure: RECO
Mk5 errors, rebooted many times.

Westerbork (Geert Kuper at Sun Nov 06 18:55:42 2011) Success:
Single dish requested: Only RT7 in TADUMax.

Torun (Magda Kunert-Bajraszewska at Mon Nov 07 10:01:56 2011) Success:
BBC6 was unlocked during this experiment. No other known problems.

Effelsberg (Uwe Bach at Mon Nov 07 13:15:31 2011) Success: LATE
Started 5min late. Stopped at 2:06 because source was setting and below elevation limit of 8 deg.

Onsala (Michael Lindqvist at Tue Nov 08 08:14:03 2011) Success:
RFI, otherwise no known problems.

Effelsberg (Uwe Bach at Tue Nov 08 08:23:47 2011) Success:
No known problems. The first comment about late start and early stop belongs to re003a.

Arecibo (Tapasi Ghosh at Tue Nov 08 20:23:51 2011) Minor failures: POIN
Due to unexpected delay in the communication system between the FS computer and the telescope's drive system, the phase referencing got hampered. A detailed list of the times when telescope was tracking the sources will be sent to the vlbeer area later. Later, it was found to be due to recently implemented changes in the internal network. connections.